OpenStudy (anonymous):

make x the subject of the formula in y=x+4/x-3

OpenStudy (yttrium):

You can multiply the whole equation by x-3 leading to y(x-3) = x+4 @prasheelasingh do you know what to do next?

OpenStudy (yttrium):

I hope you understand it with this. First, let's isolate the denominator x-3 by multiplying the whole equation by x-3 \[(x-3) [y = \frac{ x+4 }{ x-3 }]\] Therefore, \[y(x-3) =x+4 \] Let distribute y in the left side \[xy-3y = x+4\] To make x as the subject of the equation, we need to isolate all terms having x. Hence \[xy -x = 4+3y\] Then factor out x and solve for x. \[x (y-1) = 3y+4\] Therefore, \[x = \frac{ 3y+4 }{ y-1 }\]

OpenStudy (anonymous):

thanks...

OpenStudy (yttrium):

Do you understand?

OpenStudy (anonymous):

yes..